OpenStudy (anonymous):

Orlando lives in South Dakota and pays 4% in sales tax. He bought a refrigerator, and the amount he paid after sales tax was applied was $2749.87. What was the approximate cost of the refrigerator before sales tax was applied? A.$2644.11 B.$2738.91 C.$2760.87 D.$2859.86

OpenStudy (ganpat):

let cost of refrigerator be x So x + 4 % of x = 2749.87... Find x ... can u do that ??

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

Is it A?

OpenStudy (ganpat):

have u tried ??

OpenStudy (anonymous):

yup

OpenStudy (ganpat):

x + 4x/100 = 2749.87 104x / 100 = 2749.87 x = 2644.105..

OpenStudy (ganpat):

perfect !! :)

OpenStudy (anonymous):

another !!!! Gail lives in Kentucky and pays 6% in sales tax. She just bought $165 in groceries, but $80 worth of those groceries were nontaxable. What is the total amount that Gail paid for the groceries, including sales tax? A.$170.10 B.$165.99 C.$165.51 D.$174.90

OpenStudy (ganpat):

165 $ total groceries, 80 $ is tax free.. so taxable grocery = 165 -80..= 85 apply 6 % of tax on 85..

OpenStudy (anonymous):

so A?

OpenStudy (ganpat):

80 + 85 + 6% 0f 85 = 170.1 .. yup... u r good man..